Twenty students were randomly assigned to one of four exercise programs (5 students in each program), and the mean weight losses for the four programs were compared using one-way analysis of variance.The mean square error (MS Error) was 40.The mean weight loss for the first program was 1
x= 5 pounds.

39.What is the standard error of 1
x, the sample mean weight loss in the first program?
A.0.32
B.1.41
C.2.83
D.None of the above
KEY: C

How we find C(2.83)?
Answer #1

here as we know that MSE =40

therefore standard error of the sample mean weight loss in the first program =sqrt(MSE*(1/n))

=sqrt(40*(1/5))= 2.83